OpenStudy (anonymous):

It took the Gibson family 2 days to travel 925 miles to their vacation home. They traveled 75 miles more on the first day than on the second. How many miles did they travel each day?

OpenStudy (anonymous):

x+x+75=925 2x=925-75 2x=850 x=850/2 x=425 miles the second day. 425+75=500 for the first day. Proof 400+525=925 925=925

OpenStudy (anonymous):

hope you undrerstand it

OpenStudy (anonymous):

dont forget to give me a medal if I deserve one